A small ball of mass m = 5 kg is suspended from 

a string of length L = 5 m. The ball revolves 

with constant speed v in a horizontal circle of 

radius r = 2 m. Find an expression for v and a.


1
Expert's answer
2021-04-28T07:23:39-0400

Answer

By applying equilibrium in vertical

"Tcos\theta=mg"



And

"Tsin\theta=\frac{mv^2}{r}\\mgtan\theta=\frac{mv^2}{r}\\v=\sqrt{rgtan\theta}"

This is the required equation of velocity